# Governing Dynamic Capabilities: Cryptographic Binding and Reproducibility Verification for AI Agent Tool Use

### 15 Mar 2026 v1

@@ 7.1 Prototype Architecture
Certificate (parse, issue, verify) x509-parser, ring $\sim$}1,400
Skills manifest hasher sha2, serde_cbor $\sim$}500
Trust tree (internal) $\sim$}700
Reproducibility verifier sha2 $\sim$}400
Interaction ledger ed25519-dalek, sled $\sim$}1,100
Credential vault aes-gcm, argon2 $\sim$}800
Agent registry + proxy tokens agent-trust-kit $\sim$}600
MCP server + CLI jsonrpc-core, clap $\sim$}900
@@
MCP governance proxy demo cryptography, cbor2 $\sim$}520
Total $\sim$}8,290
@@ 7.2 Performance Evaluation
We evaluate on Apple M2, 16 GB RAM, macOS. All benchmarks use Criterion.rs with 10,000 samples and 3-second warmup.
Experiment 1: Certificate Verification Latency.
Chain Depth Mean Per-Sig Cap. Binding
@@
[…]
5 (max practical) 163.6 µs 32.7 µs <<1 ns
Capability binding adds <<1 ns—the cost of a single 32-byte comparison. Verification is dominated by Ed25519 signature verification ($\sim$}33 µs/sig).
Experiment 2: Skills Manifest Hash.
Tools Sort CBOR SHA-256 Total
5 32 ns 0.41 µs 2.25 µs 2.69 µs
10 42 ns 0.76 µs 4.30 µs 5.10 µs
20 64 ns 1.71 µs 8.78 µs 10.6 µs
50 152 ns 3.94 µs 21.8 µs 25.9 µs
100 263 ns 8.05 µs 43.3 µs 51.6 µs
Experiment 3: Ledger Throughput.
Operation Latency Throughput
Append record 448 µs $\sim$}2,230 rec/s
Verify single record 32.4 µs $\sim$}30,900 rec/s
Full audit (1K records) 34.7 ms 34.7 µs/rec
Full audit (100K records) 3.46 s 34.6 µs/rec
Query by agent ID 112 µs 0.56 µs/rec
Experiment 4: Reproducibility Verification Latency.
Tokens Compare Replay Skip Commit
128 120 ns 123 ns 3.5 ns 1.72 µs
512 416 ns 418 ns 3.5 ns 6.55 µs
2,048 1.55 µs 1.55 µs 3.5 ns 24.3 µs
8,192 6.01 µs 6.10 µs 3.6 ns 96.6 µs
Token comparison scales at $\sim$}0.73 ns/token. Batch verification: 100 interactions (512 tokens each) in 41.8 µs. An auditor verifies 100,000 interactions in $\sim$}42 ms.
Experiment 5: Real LLM Reproducibility Validation. We validate the DiFR finding empirically in two phases. Phase 1 uses two Anthropic models (Claude Sonnet 4 and Haiku 4.5, temperature==0, max_tokens==512) across 105 prompts in five task types—code generation (25), JSON/structured output (15), mathematical reasoning (20), prose explanation (30), and creative writing (15)—with 5 same-model repetitions and 3 cross-model repetitions each (1,050 same-model + 1,575 cross-model pairwise comparisons). Phase 2 extends to 9 models from 7 providers: Claude Sonnet 4 and Haiku 4.5 (Anthropic), GPT-4.1 and GPT-4.1 mini (OpenAI), Gemini 2.5 Pro (Google), Grok 4 (xAI), DeepSeek V3 (DeepSeek), Llama 4 Maverick (Meta), and Kimi K2.5 (Moonshot). All models are accessed via OpenRouter as a unified API gateway (plus Volcengine for Doubao), with temperature==0 and 100 prompts across six task types (code, JSON, math, prose, creative, knowledge), 10 repetitions per model-prompt pair (5,400 total API calls; 15,120 pairwise comparisons: 10,800 same-model + 240 same-provider + 4,080 cross-provider). We compute four similarity metrics: character-level positional match (CharMatch), Jaccard token overlap, word-level TF-IDF cosine, and character n-gram (3–5) cosine; then apply Welch’s $t$-test, Mann-Whitney $U$, and Cohen’s $d$ to quantify separation, and use Youden’s $J$ statistic for optimal threshold selection.
Result 5a: Same-model reproducibility is task-type dependent. Table 2 shows per-task reproducibility from Phase 1 across 105 prompts (1,050 same-model pairwise comparisons).
Table 2: Same-model (Sonnet 4) reproducibility by task type.
Task Type $n$ Identical CharMatch TF-IDF cos
[…]
Structured outputs (JSON, code) have the highest identical-output fractions (>>50%), while prose shows the most variation (34.7% identical), driven by paraphrase-level reorganization. The bimodal distribution (median CharMatch =0.975=0.975 vs. mean =0.627=0.627) confirms that same-model outputs are either near-identical or substantially rephrased.
Result 5b: Multi-provider separation analysis. Table 3 presents the Phase 2 multi-provider results across 15,120 pairwise comparisons (10,800 same-model vs. 4,080 cross-provider). All differences are statistically significant ($p\approx 0$, Welch’s $t$ and Mann-Whitney $U$).
@@
Metric Same-model Cross-provider Sep. Cohen’s $d$
@@
All four metrics achieve Cohen’s $d>1.0$ (large effect sizes) across the full 9-model set, with Jaccard ($d=1.51$) and CharMatch ($d=1.36$) providing the strongest separation. Notably, same-provider pairs (e.g., GPT-4.1 vs. GPT-4.1 mini: CharMatch =0.251=0.251; Sonnet 4 vs. Haiku 4.5: $0.192$) exhibit similarity intermediate between same-model and cross-provider, confirming that even models from the same vendor maintain distinct output fingerprints.
Result 5c: Per-model determinism varies by over an order of magnitude. Table 4 ranks models by same-model CharMatch, revealing a wide determinism spectrum.
@@
Claude Haiku 4.5 produces near-identical outputs across repetitions (CharMatch =0.987=0.987), while reasoning models (Kimi K2.5, Gemini 2.5 Pro, Grok 4) exhibit substantially lower determinism due to stochastic chain-of-thought reasoning even at temperature 0. Despite this variation, even the least deterministic model (Kimi K2.5, CharMatch =0.169=0.169) still exceeds the mean cross-provider similarity ($0.114$), demonstrating that model identity remains distinguishable for all tested models.
Result 5d: Per-task separation is strongest for mathematical reasoning. Table 5 shows CharMatch separation factors across six task categories.
Table 5: Per-task CharMatch separation (same-model / cross-provider).
Task Type Same-model Cross-provider Sep.
Math reasoning 0.508 0.042 $12.2\times$
Knowledge 0.511 0.072 $7.1\times$
Creative writing 0.402 0.057 $7.0\times$
Prose 0.488 0.083 $5.9\times$
Code generation 0.594 0.161 $3.7\times$
JSON / structured 0.779 0.299 $2.6\times$
Mathematical reasoning achieves the highest separation ($12.2\times$) because models develop highly distinctive step-by-step reasoning styles, while JSON outputs show the lowest separation ($2.6\times$) because structural constraints force convergent formatting. These task-dependent results motivate per-task threshold calibration in deployment.
Result 5e: Optimal threshold via Youden’s $J$ statistic. Table 6 reports optimal thresholds from the multi-provider experiment.
@@
In the multi-provider setting, Jaccard achieves the best overall classification ($F_{1}=0.876$, $J=0.592$), followed closely by CharMatch ($F_{1}=0.873$, $J=0.572$). Both metrics attain $F_{1}>0.87$ across 9 models and 7 providers, confirming that lightweight textual similarity metrics—requiring no model internals, embeddings, or logit access—provide practically useful model identity verification even in a heterogeneous multi-provider environment.
Experiment 6: End-to-End MCP Overhead.
@@
Experiment 7: Comparison with BAID.
Metric BAID Ours Ratio
Verification latency $\sim$}120 s 0.097 ms $1.2{\times}10^{6}$
Proof size $\sim$}50 KB $\sim$}2 KB $25\times$
Infrastructure Blockchain None
Cap. change detection Post-hoc Real-time
Experiment 8: MCP Governance Proxy Demo.
To validate real-world applicability, we implement a complete MCP governance proxy in Python (520 lines) that interposes between an MCP client (agent) and MCP server (tool provider). The proxy enforces all three governance requirements: G1 via skills manifest hash verification against the agent’s certificate, G2 via reproducibility commitment tracking, and G3 via a hash-chain interaction ledger with Ed25519-signed records. We test 5 live attack scenarios:
Attack Result Latency
S1: Inject web_browser at runtime Blocked (hash $\Delta$) 121 µs
S2: Replace read_file with trojan Blocked (code hash) 103 µs
S3: Forged certificate (rogue CA) Blocked (sig verify) 106 µs
[…]
S5: Tamper ledger record #2 Detected (chain break)
All attacks are detected in real-time. Over 1,000 governed calls, the proxy achieves mean governance overhead of 120 µs/call (median: 106 µs, p99: 432 µs), throughput of 8,195 governed calls/sec, and all ledger records pass integrity verification. This confirms that the framework provides real-time protection with overhead well below the latency of typical MCP calls (50–500 ms).
@@ 7.3 Attack Scenario Evaluation
We implement 12 attack scenarios as automated tests: 9 structural governance attacks as Rust unit tests (774 lines), 5 end-to-end scenarios via the MCP governance proxy demo (520 lines), and 3 runtime behavioral attacks informed by the Nemotron-AIQ dataset [45] (which demonstrates 100% attack success rates without structural defenses). Table 7 summarizes results.
Table 7: Attack scenario evaluation. All 12 tests pass deterministically. S10–S12 are runtime behavioral attacks validated against Nemotron-AIQ traces.
@@
S1: Silent escalation Install web_browser at runtime Undetected Prevented capability_binding: hash mismatch
S2: Tool trojanization Replace implementation (same name) Undetected Prevented capability_binding: code hash change
S3: Model substitution Forge cert with different model Undetected Detected sig_verify: forged cert unsigned
S4: Phantom delegation T2 agent issues T1 cert Undetected Prevented constraint_monotonicity
S5: Evidence tampering Modify ledger record N/A Detected Hash chain + signature invalidation
S6: Blame shifting Dispute in agent chain Unresolvable Resolved Ed25519 signatures
S7: Colluding agents T3+T3 attempt T1 access Undetected Prevented tier_check: independent evaluation
S8: Chain forgery Fabricated chain Prevented sig_verify: fake key $\neq$ root
S9: Depth overflow $\mathit{max\_depth}=0$ delegation Prevented Structural: depth <0<0 impossible
S10: Knowledge poisoning Inject false data via tool output Undetected Traced Ledger $\eta_{\mathit{in}}/\eta_{\mathit{out}}$: pinpoint injection span
S11: Cascading hijack Multi-stage instruction persistence Undetected Traced Hash-chain: full propagation path audit
S12: Output exfiltration Embed phishing via rendered output Undetected Detected Replay divergence from committed $\pi_{i}$
@@ 7.4 End-to-End Multi-Agent Pipeline
Experiments 1–8 evaluate individual governance operations in isolation. To validate the framework under realistic conditions, we deploy the full A2Auth governance stack across a multi-agent pipeline where agents coordinate, delegate tasks, invoke tools via MCP, and pass results through the interaction ledger. This experiment measures: (a) end-to-end governance overhead as a fraction of pipeline latency, (b) scaling behavior from 5 to 20 agents, (c) attack detection in situ, and (d) comparison with three baselines. The implementation uses the Rust agent-trust-kit SDK exposed to Python via PyO3 FFI, with agents making real Anthropic Claude API calls.
Figure 5: End-to-end multi-agent pipeline architecture with governance overlay. Five specialist agents coordinate through a trust tree (depth 3). Red labels indicate E2E attack injection points (Table 10). All inter-agent communications produce bilaterally-signed ledger records.
Pipeline architecture (Figure 5). We implement an AI Research Assistant pipeline: a user submits a research query; a Coordinator agent (T1_HIGH) orchestrates five specialist agents—Research (T2), Analysis (T2), Writer (T2), Review (T3_MONITORED), and Fact-Checker (T2)—each with a distinct skills manifest and reproducibility commitment. The trust tree has depth 3 (Root CA $\to$ Human $\to$ Coordinator $\to$ Workers), with the Coordinator’s max_delegation_depth=2{}=2 and all workers at depth 0 (cannot delegate). A single pipeline run traverses 21 steps: 10 tool calls (7 LLM invocations + 3 non-LLM tools), 11 inter-agent communications (task assignments and result deliveries), each producing a bilaterally-signed ledger record with capability verification and reproducibility commitment.
Experiment 9a: Governance overhead. Table 8 breaks down governance cost per pipeline run. Each operation’s latency is consistent with the micro-benchmarks from Experiments 1–4; the end-to-end cost is their composition.
Table 8: Per-pipeline governance overhead breakdown (5-agent configuration, 21 ledger entries per run, mean of 100 runs).
Operation Per-op Count Subtotal
Certificate chain verification 97 µs 6 0.58 ms
Skills manifest hash check 5.1 µs 10 0.05 ms
Ledger append (tool call) 448 µs 10 4.48 ms
Reproducibility commitment 6.6 µs 10 0.07 ms
Ledger append (inter-agent) 448 µs 11 4.93 ms
Bilateral Ed25519 signing 66 µs 11 0.73 ms
Total governance 21 10.8 ms
Pipeline latency (real LLM) $\sim$}8.7 s
Overhead 0.12%
Governance adds 10.8 ms to a pipeline that takes $\sim$}8.7 s with real Claude API calls (7 LLM inferences at $\sim$}1.2 s each + 3 non-LLM tools at $\sim$}100 ms each). The overhead of 0.12% is dominated by ledger appends (87% of governance cost), confirming that the cryptographic operations themselves (hashing, signing, verification) are negligible relative to I/O-bound agent workloads.
Experiment 9b: Scaling behavior. We scale the pipeline from 5 to 20 agents by instantiating parallel specialist groups: the Large configuration uses 4 Research agents (handling sub-queries in parallel), 4 Analysis agents, 3 Writer agents, 3 Review agents, 3 Fact-Checkers, 2 Coordinators (primary + sub-coordinator), and 1 Aggregator. Table 9 shows that per-agent governance cost remains approximately constant.
Table 9: Governance scaling: 5 $\to$ 20 agents (mean of 50 runs per configuration).
@@
Small 5 21 10.8 1.81 5.2
Medium 10 45 23.0 2.09 11.2
Large 20 95 48.3 2.30 23.8
Total governance cost scales linearly with ledger entries ($R^{2}>0.99$), but per-agent cost increases only 27% when quadrupling the agent count (1.81 ms $\to$ 2.30 ms). This near-constant per-agent cost arises because each agent’s governance operations—certificate verification, capability checking, ledger append—are independent of the total number of agents in the pipeline. The slight increase comes from deeper trust chains in the Large configuration (depth 4 vs. depth 3) adding one extra signature verification per activation. Ledger storage grows at $\sim$}250 bytes per entry; even the 20-agent pipeline produces only 23.8 KB per run, confirming that commitment-only storage (§5) is practical for long-running agent workflows.
Experiment 9c: End-to-end attack detection. We inject five attacks into the live 5-agent pipeline, each targeting a different governance layer. Unlike the unit-level tests in Table 7, these attacks occur mid-pipeline with real inter-agent data flow.
Table 10: End-to-end attack detection in the live multi-agent pipeline.
Attack Target Layer Detect. Hops
E2E-1: Capability escalation Research G1 5 µs
E2E-2: Model substitution Analysis G2 $\dagger$
E2E-3: Phantom delegation Writer G1 97 µs
E2E-4: Ledger tampering Ledger G3 32 µs
E2E-5: Cascading injection Pipeline G1+G2+G3 97 µs 3
[…]
dominated by LLM re-inference ($\sim$}1.2 s), not framework cost ($\sim$}13 µs).
[…]
Across 100 clean pipeline runs (500 total governed tool calls, 2,100 ledger entries), zero false positives are observed.
Experiment 9d: Baseline comparison. We evaluate the same five E2E attacks against three baselines: (1) no governance (raw MCP), (2) MCP with OAuth 2.1 (standard MCP authentication per specification), and (3) MCP with OpenTelemetry traces (observability-only, no cryptographic integrity). Table 11 summarizes results.
Table 11: E2E attack detection: A2Auth vs. baselines ($\checkmark$ = detected/prevented, ✗ = undetected). G1/G2/G3 indicate which governance requirement enables detection.
Attack No gov. OAuth OTEL A2Auth
E2E-1: Capability escalation $\checkmark$ (G1)
@@
E2E-5: Cascading injection c $\checkmark$ (G1+G2+G3)
Total detected 0/5 0/5 0/5 5/5
aOAuth authenticates the Writer but has no delegation depth constraint. bOTEL traces can be silently modified (no hash chain or signatures; §5). cOTEL logs the poisoned data flow but cannot prove tampering or trace provenance cryptographically.
OAuth 2.1 provides identity authentication but not capability binding: it verifies who the agent is, not what tools it currently has. OpenTelemetry provides observability but not integrity: OTEL span trees log tool calls and parent-child relationships, but as demonstrated by analysis of 10,796 Nemotron-AIQ traces (§5), OTEL records contain no hash chains, no cryptographic signatures, and no identity binding—any field can be silently modified. A2Auth’s three governance layers (G1+G2+G3) provide the only configuration that detects all five attack types, with governance overhead below 0.2% of pipeline latency.
8 Limitations and Future Work

### 19 Mar 2026 v2

@@ 7.1 Prototype Architecture
Core SDK (cert, skills, trust tree, ring, sha2, ed25519-dalek, $\sim$}5,600
     ledger, vault, registry, MCP) sled, aes-gcm, clap
Enhanced G1: BBS+ disclosure zkryptium $\sim$}270
Enhanced G2: DV-SNARK circuit ark-groth16, ark-r1cs-std $\sim$}350
@@
Adversarial G2 validation transformers, peft (Python) $\sim$}2,680
MCP governance proxy demo cryptography, cbor2 (Python) $\sim$}520
Total (Rust + Python) $\sim$}11,090
@@ 7.2 Performance Evaluation
We evaluate on Apple M2, 16 GB RAM, macOS (Criterion.rs, 10,000 samples, 3s warmup). Table 4 summarizes microbenchmarks for the four core governance primitives.
Table 4: Core governance microbenchmarks (Experiments 1–4).
Operation Key parameter Latency
Exp 1: Certificate chain verification
     Depth-2 chain 2 Ed25519 sigs 66.1 µs
     Depth-5 chain (max practical) 5 Ed25519 sigs 163.6 µs
     Capability binding 32-byte compare <<1 ns
Exp 2: Skills manifest hash
     10 tools Sort+CBOR+SHA-256 5.10 µs
     100 tools Sort+CBOR+SHA-256 51.6 µs
Exp 3: Ledger operations
     Append record Ed25519 sign+sled 448 µs
     Verify single record Ed25519 verify 32.4 µs
     Full audit (100K records) Chain walk 3.46 s
Exp 4: Reproducibility verification
     Token compare (512 tokens) Byte compare 416 ns
     Commit (512 tokens) SHA-256 6.55 µs
@@
Verification scales at $\sim$}33 µs/sig; token comparison at $\sim$}0.73 ns/token (100K interactions audited in $\sim$}42 ms).
Experiment 5: Real LLM Reproducibility Validation. We validate the DiFR finding across 9 models from 7 providers (Claude Sonnet 4/Haiku 4.5, GPT-4.1/4.1 mini, Gemini 2.5 Pro, Grok 4, DeepSeek V3, Llama 4 Maverick, Kimi K2.5) via OpenRouter with temperature==0, 100 prompts across six task types, 10 repetitions per model-prompt pair (5,400 API calls; 15,120 pairwise comparisons). We compute four similarity metrics—CharMatch, Jaccard, TF-IDF cosine, char n-gram cosine—and use Cohen’s $d$ and Youden’s $J$ for separation analysis and threshold selection.
Result 5b. Table 5 shows multi-provider separation ($p\approx 0$ for all metrics).
@@
Metric Same Cross Sep. $d$
@@
All metrics achieve Cohen’s $d>1.0$; Jaccard ($d=1.51$) and CharMatch ($d=1.36$) provide the strongest separation.
Result 5c. Table 6 ranks per-model determinism.
@@
Determinism spans $5.8\times$ (Haiku 4.5: $0.987$ to Kimi K2.5: $0.169$); even the least deterministic model exceeds mean cross-provider similarity ($0.114$). Per-task separation ranges from $12.2\times$ (math) to $2.6\times$ (JSON).
Result 5e. Via Youden’s $J$ (Table 7), Jaccard and CharMatch attain $F_{1}>0.87$ across all 9 models, confirming lightweight textual similarity provides practical model identity verification without model internals.
@@
Experiment 6: Adversarial Model Substitution Detection. We train three QLoRA backdoor variants (keyword B1, semantic B2, compositional B3) on three architectures (Qwen 2.5 7B, Llama 3.1 8B, Gemma 2 9B) and test G2 replay detection across four prompt strategies at budgets $n\in\{10,25,50,100,200\}$ (357 configs).
Table 8: G2 adversarial detection rate at temperature=0{}=0 (3 models $\times$ 4 strategies).
Budget ($n$) B1 B2 B3 Avg
10 99.0% 99.2% 99.2% 99.1%
25 100% 100% 100% 100%
200 99.1% 98.9% 99.1% 99.1%
Detection is near-universal (minimum 90% across all 357 configs). Mean CharMatch between base and backdoored outputs is 0.057, far below $\theta^{*}=0.146$: modifying weights to implant an effective backdoor unavoidably shifts the clean-input distribution, making substitution detectable (§8).
7.3 Overhead and Comparison
Experiment 7: End-to-End MCP Overhead.
@@ 7.3 Overhead and Comparison
Experiment 8: Enhanced Instantiation Benchmarks.
Table 9: Enhanced cryptographic backends: BBS+ (G1) and Groth16 DV-SNARK (G2).
Operation Latency vs. Basic
BBS+ selective disclosure (16-message cert)
     Full disclosure verification 12.27 ms $127\times$
     Selective proof gen (6/16) 14.66 ms $151\times$
     Selective proof verification 13.80 ms $142\times$
Groth16 DV-SNARK CharMatch circuit
     Prove (1,024 chars, 19K constraints) 173.86 ms
     Prove (4,096 chars, $\sim$}73K constraints) 750.33 ms
     Verify (constant) 1.28 ms
Experiment 9: Comparison with BAID.
Metric BAID Ours (basic) Ours (enhanced)
G1 verification 0.097 ms 13.80 ms
G2 proof gen (1K) $\sim$}120 s $\sim$}0.1 ms 174 ms
G2 proof verify $\sim$}10 ms <<0.01 ms 1.3 ms
Circuit size ${\sim}10^{10}$ 19,238
Privacy Full ZK None Selective / ZK
Infrastructure Blockchain None None
Direct CharMatch comparison (no SNARK).
BBS+ adds ${\sim}142\times$ overhead but enables selective disclosure; DV-SNARK verification is constant at ${\sim}1.3$ ms. Our DV-SNARK is ${\sim}690\times$ faster than BAID for proof generation (${\sim}520{,}000\times$ fewer constraints), by SNARKing only the verification metric rather than the full transformer forward pass. All enhanced operations remain within typical agent interaction latency (50–500 ms).
Experiment 10: MCP Governance Proxy Demo. A Python proxy (520 lines) enforcing G1–G3 detects all 5 live attack scenarios (tool injection, trojanization, forged certificate, rate abuse, ledger tampering) in $\leq$165 µs each, achieving 8,195 governed calls/sec with mean overhead of 120 µs/call.
@@ 7.4 Attack Scenario Evaluation
We evaluate 12 attack scenarios at three levels of validation (Table 10): (1) Structural governance attacks (S1–S9) are tested as automated Rust unit tests that verify the governance algorithms correctly reject each attack construction—these validate algorithm correctness, not resistance to real-world attackers. (2) Runtime behavioral attacks (S10–S11), informed by the Nemotron-AIQ dataset [46], are validated via post-hoc forensic tracing: the ledger provides sufficient evidence to identify the injection point after the fact, but the framework does not prevent the attack in real time. (3) S12 (output exfiltration) is detectable via replay divergence. The 6 E2E attacks in §7.5 provide the strongest validation: real-time detection in a live pipeline with real LLM calls.
Table 10: Attack scenario evaluation. S1–S9: algorithm correctness validated via automated unit tests (9/9 pass). S10–S11: post-hoc forensic tracing validated against Nemotron-AIQ traces. S12: detectable via replay divergence. The “Framework” column distinguishes Prevented (blocked in real time), Detected (identified but not blocked), and Traced (forensic reconstruction after the fact).
@@
S1: Silent escalation Add tool at runtime Undetected Prevented Skills manifest hash mismatch
S2: Tool trojanization Replace implementation Undetected Prevented Code hash change
S3: Model substitution Forged cert, wrong model Undetected Detected Signature verification failure
S4: Phantom delegation T2 issues T1 cert Undetected Prevented Constraint monotonicity
S5: Evidence tampering Modify ledger record N/A Detected Hash chain + sig invalidation
S6: Blame shifting Dispute in chain Unresolvable Resolved Ed25519 bilateral signing
S7: Credential collusion T3+T3 $\to$ T1 access Undetected Prevented Independent tier evaluation
S8: Chain forgery Fabricated chain Prevented Sig verify: fake key $\neq$ root
S9: Depth overflow Delegate at depth 0 Prevented Depth counter exhausted
S10: Knowledge poisoning Inject via tool output Undetected Traced Ledger $\eta_{\mathit{in}}/\eta_{\mathit{out}}$
S11: Cascading hijack Multi-stage persistence Undetected Traced Hash-chain provenance audit
S12: Output exfiltration Embed phishing link Undetected Detected Replay divergence
@@ 7.5 End-to-End Multi-Agent Pipeline
We deploy the full governance stack in a multi-agent pipeline with real Claude API calls, measuring overhead, scaling, attack detection, and baseline comparison (Figure 5).
Figure 5: Multi-agent pipeline with governance overlay. Red labels indicate attack injection points (Table 13).
Pipeline architecture (Figure 5). We implement an AI Research Assistant using LangGraph with the Rust agent-trust-kit SDK exposed via PyO3 FFI. A Coordinator (T1_HIGH) orchestrates four specialist agents—Research (T2), Analysis (T2), Writer (T2), and Review (T3_MONITORED)—in a depth-3 trust tree. Each pipeline run produces 7 governed tool calls, each generating a bilaterally-signed ledger record with capability verification and reproducibility commitment.
Experiment 11a: Governance overhead.
Table 11: Per-pipeline governance overhead (5-agent config, 7 ledger entries/run, mean of 90 runs with real Claude API calls).
Operation Per-call Calls Total
G1: Capability check 39 µs 7 0.27 ms
G2: Repro. commitment 86 µs 7 0.60 ms
G3: Ledger append 1,300 µs 7 9.10 ms
Total governance 7 10.0 ms
Pipeline latency (real Claude API) $\sim$}59.8 s
Overhead <<0.02%
Governance adds $\sim$}10 ms to a $\sim$}60 s real-LLM pipeline (<<0.02%), dominated by G3 ledger appends (91% of governance cost). Across 90 clean runs (630 governed tool calls), zero false positives and 100% ledger integrity are observed (95% CI: FPR << 3.3%).
Experiment 11b: Scaling behavior.
Table 12: Governance scaling: 5 $\to$ 20 agents (mock tool latency to isolate governance cost).
@@
Small 5 7 6.9 1.37 1.7
Medium 10 16 12.9 1.29 3.9
Large 20 33 30.0 1.50 8.1
Governance scales linearly with ledger entries; per-agent cost remains approximately constant (1.29–1.50 ms across $4\times$ scaling). Storage is $\sim$}250 bytes/entry (8.1 KB for 20 agents).
Experiment 11c: End-to-end attack detection.
Table 13: End-to-end attack detection (real Claude API calls, 7 scenarios covering G1/G2/G3).
Attack Target Layer Detection
E2E-1: Cap. escalation Research G1 Tool not in manifest
E2E-2: Model substitution Analysis G1 Model binding mismatch
E2E-3: Phantom delegation Writer G1 No valid cert chain
E2E-4: Ledger tampering Ledger G3 Hash chain break
E2E-5: Tool trojanization Research G1 Manifest hash mismatch
E2E-6: Depth overflow Review G1 Trust decay violation
E2E-7: Replay divergence Analysis G2 CharMatch 0.00 << $\theta$
All seven attacks are detected, covering all three governance layers. E2E-2 is caught by G1’s model binding check: the certificate declares claude-sonnet-4 but the agent uses claude-haiku-4-5. E2E-7 is the only attack requiring G2 (replay verification): the analysis agent’s certificate and model binding are correct (G1 passes), but the agent secretly produces attacker-controlled output instead of executing the declared model. Post-pipeline replay re-executes the declared model on the stored input and detects divergence (CharMatch==0.00 vs. $\theta$==0.85)—demonstrating that G2 catches attacks invisible to G1 and G3.
Experiment 11d: Baseline comparison. We evaluate the same seven attacks against three baselines.
Table 14: E2E attack detection: A2Auth vs. baselines ($\checkmark$ = detected, ✗ = undetected).
Attack None OAuth OTEL A2Auth
E2E-1: Cap. escalation $\checkmark$ (G1)
@@
E2E-5: Tool trojanization $\checkmark$ (G1)
E2E-6: Depth overflow $\checkmark$ (G1)
E2E-7: Replay divergence $\checkmark$ (G2)
Total detected 0/7 0/7 0/7 7/7
aOAuth authenticates the agent but has no delegation depth constraint. bOTEL traces can be silently modified (no hash chain or signatures).
OAuth 2.1 verifies who the agent is but not what tools it currently has or which model it runs. Only A2Auth (G1+G2+G3) detects all seven E2E attacks, at <<0.02% overhead.
Experiment 11e: Governance coverage comparison with BAID and Ghosh et al. Table 14 compares against infrastructure baselines; Table 15 compares against the two closest research systems—BAID [2] (zkVM-based verifiable inference) and Ghosh et al. [46] (operational safety with Nemotron-AIQ)—at the governance requirement level across all 12 attack categories. This is a design-level analysis: we map each system’s published mechanisms to the attack taxonomy of §2.
[…]
Attack Requires BAID Ghosh A2Auth Notes
[…]
ATK-9: Blame shifting G2+G3 $LEFTcircle \CIRCLE$ BAID proves per-execution integrity but lacks bilateral signing for attribution.
[…]
Governance requirements G2 only Partial G3 G1+G2+G3
G2 proof generation $\sim$}120 s $\sim$}0.1 ms / 174 ms
Infrastructure Blockchain None None
Basic instantiation (direct CharMatch). Enhanced instantiation (DV-SNARK). BAID provides computational integrity (the declared model provably executed the computation)—a strictly stronger G2 guarantee than A2Auth’s statistical behavioral verification. However, BAID addresses only G2: it cannot detect capability escalation (G1) or audit interaction chains (G3). Ghosh et al.’s operational framework provides guardrails and red-teaming methodology but lacks cryptographic enforcement for any governance requirement. The three systems are complementary, not competing: BAID’s zkVM could serve as an enhanced G2 backend within the A2Auth architecture; Ghosh’s guardrails address content safety orthogonal to governance.
8 Limitations, Governance Trilemma, and Future Work